OpenStudy (anonymous):

Given f(x)=2x-2/4, solve for f^-1(3).

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

translate as "solve for \(x\), \[\frac{2x-2}{4}=3\]

OpenStudy (anonymous):

btw what idiot writes \(f(x)=\frac{2x-2}{4}\) when it is just \(f(x)=\frac{x-1}{2}\) ? no matter you can solve in any case

OpenStudy (anonymous):

Lool math is hard, man.

OpenStudy (anonymous):

\[\frac{2x-2}{4}=3\\ 2x-2=12\] etc

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Yes! Thanks again!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

(I have one more)